The 1.50-cm-long second hand on a watch rotates smoothly.

A)What is its angular velocity?

b)What is the speed of the tip of the hand?
Answer #1

a) angular velocity = 2pi/60 = 0.104 rad/s

b) from

    v = rW = 1.50*10^-2* 0.104 = 1.57 x 10^-3 m/s
Answer #2

angular velocity \omega =\frac{\Theta }{t}

\omega =\frac{2\prod }{60}=0.104 rad/s

b) linear velocity v=2\prod r/t

                                   =0.00157 m/s